Required information The following information applies to the questions displayed below.] The Shirt Shop had the following transactions for T-shirts for 2018, its first year of operations: Jan. 20 Purchased Apr. 21 Purchased July 25 Purchased Sept. 19 Purchased 90 units 3-$1,170 110 units 4-440 230 units 61,380 90 units720 During the year, The Shirt Shop sold 620 T-shirts for $13 each. Compute the amount of ending inventory The Shirt Shop would report on the balance sheet, assuming the FIFO cost flow assumption FIFO Ending inventory...
Answer #1

As per FIFO;ending inventory =(110 units@$6)+(90 units@$8)

=$1380

NOTE:

As per FIFO;goods purchased first are sold off first.Hence 620 units sold would consist of 390 units of Jan 20;110 units if Apr 21 and the balance (620-390-110)=120 units of July 25

Hence ending inventory would consist of =(230-120)=110 units of July 25 and 90 units of Sept 19.